# S2R-HDR: A Large-Scale Rendered Dataset for HDR Fusion

### v1

@@ Abstract
The generalization of learning-based high dynamic range (HDR) fusion is often limited by the availability of training data, as collecting large-scale HDR images from dynamic scenes is both costly and technically challenging. To address these challenges, we propose S2R-HDR, the first large-scale high-quality synthetic dataset for HDR fusion, with 24,000 HDR samples. Using Unreal Engine 5, we design a diverse set of realistic HDR scenes that encompass various dynamic elements, motion types, high dynamic range scenes, and lighting. Additionally, we develop an efficient rendering pipeline to generate realistic HDR images. To further mitigate the domain gap between synthetic and real-world data, we introduce S2R-Adapter, a domain adaptation designed to bridge this gap and enhance the generalization ability of models. Experimental results on real-world datasets demonstrate that our approach achieves state-of-the-art HDR reconstruction performance. Dataset and code will be available at https://openimaginglab.github.io/S2R-HDR.
Figure 1: Comparing HDR fusion models [21] trained on our S2R-HDR dataset, with the proposed domain adapter S2R-Adapter, with the same model trained on previous SCT [41] and Challenge123 [21] datasets. Results show our dataset and training scheme can reduce ghosting artifacts under large motion (left) and recover very high dynamic range scenes, such as direct sunlight (right).
@@ 1 Introduction
High dynamic range (HDR) reconstruction plays a crucial role in various real-world applications, such as computational photography, visual perception, and autonomous driving. Despite notable advancements in HDR image reconstruction techniques [49, 19, 27, 41, 21] in recent years, models trained on small-scale datasets still face limitations in generalizing to complex scenes. Additionally, due to limited data scale, the complexity and challenges of HDR reconstruction have yet to be fully explored, particularly in scenarios involving large motion and direct sunlight, as illustrated in Fig. 1.
@@
In real-world scenarios, collecting comprehensive, high-quality large-scale HDR datasets for dynamic scenes is time-consuming, resource-intensive, and poses significant technical challenges. Uncontrollable elements such as lighting conditions, weather variations, and dynamic objects like animals and vehicles make it difficult to fully control the data acquisition process. Capturing extreme high dynamic range scenarios—such as environments with direct sunlight—poses an even greater challenge, like Fig. 1. Consequently, existing HDR datasets [19, 41, 21, 3, 37] are generally limited to artificially controlled dynamic scenes and fail to capture the diversity of real-world environments. For example, some datasets focus exclusively on human motion, overlooking other essential dynamic elements, such as animals and vehicles. Moreover, existing HDR datasets with ground truth fusion results are typically small. For instance, Kong et al. [21] built the latest dataset with 123 samples. Models trained on these small datasets are prone to overfitting, further hindering progress in HDR research.
Figure 2: The distribution of our S2R-HDR dataset and real captured HDR datasets [19, 41, 21]. Following the approach outlined in [37, 12, 18], we first extract 7-dimensional features that capture key aspects of HDR, including the extent of dynamic range, intra-frame diversity, and the overall style of the HDR images. These features are then projected into a 2D space using t-SNE [45] for visualization.
@@
To address these limitations, we introduce S2R-HDR, the first large-scale high-quality HDR synthetic dataset designed for HDR fusion. S2R-HDR features several distinctive characteristics: 1) High Quality: Inspired by prior works [23, 54, 16, 5, 56], we render high-quality raw HDR data using Unreal Engine, with realistic lighting, shadow, weather, and motion effects. 2) Large Scale: The dataset contains 24,000 HDR images, around 6 times larger than existing ones. 3) Diversity: The dataset encompasses different motion types and lighting. It also covers different dynamic elements such as animals, humans, and vehicles across a variety of indoor and outdoor settings. 4) Controllable Environment: Using tools developed based on xrfeitoria [7], we can flexibly control environmental factors to create diverse data.
@@
While rendering engines can generate a large volume of high-quality synthetic data, a domain gap exists between synthetic and real data, particularly in texture distribution, as shown in Fig. 2. To address this, we propose S2R-Adapter, a plug-and-play simulation-to-real domain adaptation approach designed to bridge this gap. This approach can be applied to both labeled and unlabeled data, meaning even if the target real HDR datasets do not have the ground truth fusion result, we can still adapt to it. To achieve this, inspired by previous works [14, 53, 24], our S2R-Adapter consists of two branches: 1) A share branch manages knowledge sharing, which ensures the knowledge learned from synthetic data are not forgotten, and 2) a transfer branch facilitates knowledge transfer, which ensures the model can adapt to real input.
@@
Additionally, our training strategy can be applied to different network structures, including both CNN-based and transformer-based models. Integrating this strategy using re-reparameterization [8] incurs no extra computational overhead during inference.
@@
Experimental results on both labeled and unlabeled real datasets demonstrate that the proposed dataset and method significantly enhance the performance of HDR reconstruction models trained on synthetic data when applied to real scenes, achieving state-of-the-art results. Our study not only provides a new solution for HDR reconstruction but also presents a feasible path for generalization in fields where data acquisition is challenging.
@@ 2 Related Works
Image HDR datasets Datasets are essential for the development and evaluation of algorithms. Before the deep learning era, Sen et al. [35] and Tursun et al. [43] provided real-world HDR datasets containing 8 and 16 scenes, respectively, and Kalantari and Ramamoorthi [19] further introduced the first paired LDR-HDR dataset with 89 pairs. Prabhakar et al. [32] later expanded this to 582 LDR-HDR pairs and Tel et al. [41] collected a dataset focusing on foreground objects and larger motion variations, with 144 samples. Other datasets are also built for deghosting [37], mobile imaging [25] , or large motion [21].
Image HDR methods Deep learning has been introduced into the field of HDR reconstruction due to its remarkable performance in image processing. Early researchers designed an alignment and fusion pipeline [19, 48]. Subsequent works et al. [2, 6, 26, 51] focused on improving the alignment process by developing more advanced modules to handle motion artifacts across different exposures. Kong et al. [21] also proposed a novel efficient processing network.
Over time, several alternative pipelines HDR reconstruction have been proposed, using attention mechanisms [49], non-local blocks [50], generative adversarial network et al. [30], or multi-step fusion et al. [55]. Recently, transformer models have shown promising results in HDR fusion [38, 27]. Steven Tel et al. [41] also developed a semantic-consistent, alignment-free transformer for HDR reconstruction. At last, diffusion models are also first introduced to HDR fusion by Yan et al. [52] and Hu et al. [17] further accelerate it using a low-frequency aware model.
Sim-to-real domain adaptation Domain adaptation has been widely used to transfer models trained on synthetic data to real-world settings. To address the potential domain shifts, researchers use either adversarial approaches [10, 44] or domain randomization [42]. Recently, adapter-based domain adaptation [14, 4, 39] has been proven to be more effective. Adapters [14, 4] are a form of parameter-efficient fine-tuning (PEFT) [14, 58, 11, 15], which require fewer parameters than full retraining and help mitigate catastrophic forgetting [4, 24] in domain adaptation. Additionally, Test-Time Adaptation (TTA) [22, 1, 46, 24, 4] has been extensively explored, aiming to adapt a pre-trained model to unknown target domains during test-time, without any labeled or source domain data.
@@ 3 S2R-HDR Dataset
Previously, to create an HDR dataset with ground truth, researchers often use a beam splitter and two cameras to simultaneously capture images with two different exposures [9, 47]. The beam splitter only has two different exposures, which limits the dynamic range of the image. However, there are various high dynamic range scenarios in natural scenes, such as environments with direct sunlight. Accurately extracting tens of thousands of data samples from these scenes is a significant challenge. Previously, the largest commonly used dataset contained only 144 images, whereas ours includes 24,000 HDR images, representing a substantial leap in scale and diversity.
Moreover, capturing the ground truth often requires capturing different exposure images frame-by-frame [19, 41, 21, 3, 37] and manually controlling motion between frames, making capturing extremely time-consuming. The captured motions are often limited and unrealistic, most of them are just basic human movements. These limitations have made it difficult to scale HDR datasets both in terms of size and motion variety. Below we discuss how we solve all these challenges in our synthetic rendering dataset.
@@ 3.1 Rendering Design
Rendering high-quality HDR data presents several challenges. One challenge is that rendered images have a different distribution compared to the actual raw sensor data captured by cameras. To mitigate this difference, we made several improvements. First, by default, rendered images have a baked-in tone mapping, an irreversible process that compresses dynamic range for standard displays, making it hard to recover original HDR data. To overcome this, we design a custom UE5 rendering pipeline that modifies tone mapping and gamma correction, ensuring the output remains in linear HDR space, and stores results in floating-point formats (EXR) to prevent data quantization. This approach ensures greater accuracy and makes the rendered data more suitable for HDR-related tasks. Second, we also simulate imperfections during handheld capturing. We incorporated camera shake simulation into our camera pose control to replicate the vibrations and instabilities that occur during real-world capture. This ensures that the rendered data closely mimics real-world shooting conditions, yielding more realistic HDR data for image processing and model training.
Another challenge is to construct realistic and diversified HDR scenes, with varying motion, lighting, and environmental details. To tackle this, we design and curate a diverse range of dynamic scene materials, including common moving objects such as animals, pedestrians, and vehicles, ensuring that the scenes exhibit a high degree of dynamism and complexity, as shown in Fig. 3. Additionally, we carefully build a variety of high dynamic range scenes, encompassing both indoor and outdoor environments, various lighting conditions across different times of day, and extreme lighting scenarios. This diversity ensures that the generated HDR data simulates a broad range of real-world environments as much as possible.
In total, we rendered 1,000 sequences, each containing 24 frames, resulting in a dataset of 24,000 HDR images, all stored in EXR format at a resolution of 1920 × 1080. As demonstrated in Fig. 3, our rendered data encompasses a variety of environments and includes a broad range of motion types, showcasing a high degree of variability. Furthermore, since the data is in linear HDR format, it facilitates flexible data augmentation, enabling the easy generation of different LDR (low dynamic range) images, as shown in Fig. 4.
@@ 3.2 Statistics and Analysis
Dataset FHLP $\uparrow$ EHL $\uparrow$ SI $\uparrow$ CF $\uparrow$ stdL $\uparrow$ ALL $\uparrow$ DR $\uparrow$
Kalantari [19] 15.07 3.07 18.4 4.74 10.02 6.19 2.71
SCT [41] 12.43 2.43 18.25 3.92 9.39 5.44 2.55
Challenge123 [21] 26.91 5.19 20.47 5.19 12.73 9.88 2.36
S2R-HDR 28.02 5.47 38.02 14.96 15.16 10.53 3.86
We further analyze diversity of S2R-HDR in comparison to previous datasets [19, 41, 21]. Following the methodology of [37, 12, 18], we use seven metrics to evaluate the diversity of different datasets across three dimensions: the extent of HDR, intra-frame diversity, and overall HDR style. As shown in Tab. 1, the S2R-HDR dataset outperforms all prior datasets across these metrics. The “Extent of HDR” metric demonstrates that our dataset covers a broader range of highlights, indicating an extended highlight range. The “Intra-frame Diversity” metric suggests that our images contain more detailed information and richer content. Finally, the “Overall Style” metric reveals that S2R-HDR exhibits a significantly higher dynamic range, surpassing the performance of previous datasets. Details of seven metrics can be found in the appendix Sec. B.4.
Additionally, to visually illustrate the distribution between our dataset and existing real-world datasets [19, 41, 21], we extract seven-dimensional feature vectors for each image and apply t-SNE [45] for dimensionality reduction. As shown in Fig. 2, our S2R-HDR dataset spans a broader range in terms of data diversity.
@@ 4 Domain Adaption
Figure 5: Structure of S2R-Adapter and t-SNE visualization of feature representations from different branches.
With all the careful design proposed in the previous section, there is still a noticeable gap between the synthetic S2R-HDR dataset and the real one, as shown in the t-SNE visualization in Fig. 2. Thus, it is crucial to adapt the model trained on a large-scale rendered dataset to a small-scale real one. Still, direct fine-tuning on labeled real data can lead to overfitting and knowledge forgetting [57, 20].
@@
This is inspired by recent studies [24, 4, 39], which suggest that adapters [14, 33, 4] can mitigate forgetting in high-level vision tasks. Our adapter consists of two branches: a share branch to preserve shared knowledge from the rendered dataset, and a transfer branch to learn domain-specific knowledge from the real dataset, as shown in Fig. 5(a). We chose this design because we want to utilize both the shared knowledge from S2R-HDR to address large motion and dynamic range fusion, and the domain-specific knowledge from the real dataset, like more realistic textures.
More specifically, the proposed S2R-Adapter uses a plug-and-play structure, which can be attached to any pre-trained layers performing matrix multiplication (e.g., Linear Layer, Convolution Layer). Following  [24], we use a low-rank adapter as the share branch, which can better address knowledge forgetting, and use a high-rank adapter as the transfer branch, which can better extract domain-specific knowledge. Below we introduce details of each branch.
@@
$f=W_{0}x+\alpha_{s}\times f_{s}+\alpha_{t}\times f_{t}.\vspace{-10pt}$ (1)
@@
Verification using t-SNE. To verify the effectiveness of the proposed share branch and transfer branch adapters, we visualize the distributions of the rendered and real images using t-SNE [45] in  Fig. 5 (b). From the share branch adapter, the feature distributions are consistent between the real and rendered domain, indicating that the share branch can ignore the domain difference between the real and the rendered domain, preserving the shared knowledge from forgetting. On the other hand, the transfer branch better separates the real distribution from the rendered distribution, showing its capability to model the real data distribution better and extract domain-specific knowledge in the real domain.
@@
Test-time adaptation with unlabeled data. During test-time adaptation, no labeled real domain data is available, and each sample is seen only once. Therefore, $\alpha_{s}$ and $\alpha_{t}$ cannot be learned across the real domain. Moreover, each test sample’s varying distance to the rendered domain requires adaptive scaling of transfer and shared branches. Therefore, inspired by [24, 31], we dynamically adjust the scale factors using domain shift. For larger shifts, we increase the transfer branch’s scale factor, encouraging more knowledge from the real domain. For smaller shifts, we allocate more from the shared branch, preserving rendered domain knowledge. Domain shift is measured by uncertainty, following [46, 24, 34, 31]. In our HDR Fusion task, we augment input samples $N$ times and calculate variance across $N$ outputs as the uncertainty value $\mathcal{U}(x)$. Augmentations include adjusting exposure, white balance, noise levels, and random flips. With the uncertainty value, we adaptively adjust scale factors:
$\alpha_{s}=1-\mathcal{U}(x);\quad\alpha_{t}=1+\mathcal{U}(x).\vspace{-0.3cm}$ (2)
Following previous works on test-time adaptation [53, 46], we utilize the mean-teacher framework. As shown in Fig. 6, we inject S2R-Adapters to both the teacher model $\mathcal{T}$ and the student model $\mathcal{S}$. We initialized both models with pre-trained weights on the rendered domain. Following [24], the teacher model generates uncertainty values and pseudo-labels $\widetilde{y}$ for updating the S2R-Adapters. The student model is optimized by the loss between the student output $\hat{y}$ and the pseudo-label $\widetilde{y}$. The teacher model updates via the exponential moving average (EMA) of the student model:
$\vspace{-0.1cm}\mathcal{T}^{t}=\lambda\mathcal{T}^{t-1}+(1-\lambda)\mathcal{S}^{t},\vspace{-0.1cm}$ (3)
where $t$ is the test step, $\lambda$ is set to $0.999$, following [40].

### v2

@@ Abstract
The generalization of learning-based high dynamic range (HDR) fusion is often limited by the availability of training data, as collecting large-scale HDR images from dynamic scenes is both costly and technically challenging. To address these challenges, we propose S2R-HDR, the first large-scale high-quality synthetic dataset for HDR fusion, with 24,000 HDR samples. Using Unreal Engine 5, we design a diverse set of realistic HDR scenes that encompass various dynamic elements, motion types, high dynamic range scenes, and lighting. Additionally, we develop an efficient rendering pipeline to generate realistic HDR images. To further mitigate the domain gap between synthetic and real-world data, we introduce S2R-Adapter, a domain adaptation designed to bridge this gap and enhance the generalization ability of models. Experimental results on real-world datasets demonstrate that our approach achieves state-of-the-art HDR fusion performance. Dataset and code are available at https://openimaginglab.github.io/S2R-HDR.
@@ 1 Introduction
High dynamic range (HDR) fusion plays a crucial role in various real-world applications, such as computational photography, visual perception, and autonomous driving. Despite notable advancements in HDR fusion techniques (Yan et al., 2019; Kalantari and Ramamoorthi, 2017; Liu et al., 2022; Tel et al., 2023; Kong et al., 2024) in recent years, models trained on small-scale datasets (Kalantari and Ramamoorthi, 2017; Chen et al., 2021; Kong et al., 2024; Tel et al., 2023) still face limitations in generalizing to complex scenes. Additionally, due to limited data scale, the complexity and challenges of HDR fusion have yet to be fully explored, particularly in scenarios involving large motion and direct sunlight, as illustrated in Figure  1.
@@
In real-world scenarios, collecting comprehensive, high-quality large-scale HDR datasets for dynamic scenes is time-consuming, resource-intensive, and poses significant technical challenges. Uncontrollable elements such as lighting conditions, weather variations, and dynamic objects like animals and vehicles make it difficult to fully control the data acquisition process. Capturing extreme high dynamic range scenarios—such as environments with direct sunlight—poses an even greater challenge, like Figure 1. Consequently, existing HDR datasets (Kalantari and Ramamoorthi, 2017; Tel et al., 2023; Kong et al., 2024; Chen et al., 2021; Shu et al., 2024) are generally limited to artificially controlled dynamic scenes and fail to capture the diversity of real-world environments. For example, some datasets focus exclusively on human motion, overlooking other essential dynamic elements, such as animals and vehicles. Moreover, existing HDR datasets with ground truth fusion results are typically small. For instance,  Kong et al. (2024) built the latest dataset with 123 samples. Models trained on these small datasets are prone to overfitting, limiting their performance under challenging scenarios. A larger synthetic dataset (Barua et al., 2025) has been proposed recently for the single-image LDR to HDR conversion task, which is intrinsically limited by the absence of complementary exposure information, resulting in unrecoverable details in saturated regions.
Figure 1: Comparing HDR fusion models (Kong et al., 2024) trained on our S2R-HDR dataset, with the proposed domain adapter S2R-Adapter, with the same model trained on previous SCT (Tel et al., 2023) and Challenge123 (Kong et al., 2024) datasets. Results show our dataset and training scheme can reduce ghosting artifacts under large motion (left) and recover very high dynamic range scenes, such as direct sunlight (right).
@@
To address these limitations, we introduce S2R-HDR, the first large-scale HDR synthetic dataset designed for HDR fusion. S2R-HDR features several distinctive characteristics: 1) High Quality: Inspired by prior works (Li et al., 2023; Yang et al., 2023; Hu et al., 2023; Chen et al., 2023; Yin et al., 2024), we render high-quality raw HDR data using Unreal Engine, with realistic lighting, shadow, weather, and motion effects. 2) Large Scale: The dataset contains 24,000 HDR images, around 166 times larger than typical datasets (Kalantari and Ramamoorthi, 2017; Tel et al., 2023; Kong et al., 2024). 3) Diversity: The dataset encompasses different motion types and lighting. It also covers different dynamic elements such as animals, humans, and vehicles across a variety of indoor and outdoor settings. 4) Controllable Environment: Using tools developed based on xrfeitoria (Contributors, 2023), we can flexibly control environmental factors to create diverse data.
@@
While rendering engines can generate a large volume of high-quality synthetic data, a domain gap exists between synthetic and real data, particularly in texture distribution, as discussed in Appendix A.2. To address this, we propose S2R-Adapter, a plug-and-play simulation-to-real domain adaptation approach designed to bridge this gap. This approach can be applied to both labeled and unlabeled data, meaning even if the target real HDR datasets do not have the ground truth fusion result, we can still adapt to it. To achieve this, inspired by previous works (Hu et al., 2021; Yang et al., 2024; Liu et al., 2023a), our S2R-Adapter consists of two branches: 1) A share branch manages knowledge sharing, which ensures the knowledge learned from synthetic data are not forgotten, and 2) a transfer branch facilitates knowledge transfer, which ensures the model can adapt to real input.
@@
Additionally, our training strategy can be applied to different network structures, including both CNN-based and transformer-based models. Integrating this strategy using re-reparameterization (Ding et al., 2021) incurs no extra computational overhead during inference.
@@
Experimental results on both labeled and unlabeled real datasets demonstrate that the proposed dataset and method significantly enhance the performance of HDR fusion models trained on synthetic data when applied to real scenes, achieving state-of-the-art results. Our study not only provides a new solution for HDR fusion but also presents a feasible path for generalization in fields where data acquisition is challenging.
@@ 2 Related Works
Image HDR datasets. Datasets are essential for the development and evaluation of algorithms. Before the deep learning era,  Sen et al. (2012) and  Tursun et al. (2016) provided real-world HDR datasets containing 8 and 16 scenes, respectively, and  Kalantari and Ramamoorthi (2017) further introduced the first paired LDR-HDR dataset with 89 pairs.  Prabhakar et al. (2019) later expanded this to 582 LDR-HDR pairs and  Tel et al. (2023) collected a dataset focusing on foreground objects and larger motion variations, with 144 samples. Other datasets are also built for deghosting (Shu et al., 2024), mobile imaging (Liu et al., 2023b) , or large motion (Kong et al., 2024).
Image HDR methods. Deep learning has been introduced into the field of HDR fusion due to its remarkable performance in image processing. Early researchers designed an alignment and fusion pipeline (Kalantari and Ramamoorthi, 2017; Wu et al., 2018). Subsequent works (Catley-Chandar et al., 2022; Chung and Cho, 2023; Liu et al., 2021; Yan et al., 2023a) focused on improving the alignment process by developing more advanced modules to handle motion artifacts across different exposures.  Kong et al. (2024) also proposed a novel efficient processing network.
Over time, several alternative pipelines for HDR fusion have been proposed, using attention mechanisms (Yan et al., 2019), non-local blocks (Yan et al., 2020), generative adversarial networks (Niu et al., 2021), or multi-step fusion (Ye et al., 2021). Recently, transformer models have shown promising results in HDR fusion (Song et al., 2022; Liu et al., 2022).  Tel et al. (2023) also developed a semantic-consistent, alignment-free transformer for HDR fusion. Recently, diffusion models have also been introduced to HDR fusion by  (Yan et al., 2023b; Hu et al., 2024; Chen et al., 2025), further improving the performance. Self-supervised approaches have also been introduced to HDR fusion, with the SelfHDR method (Zhang et al., 2024) utilizing a color and structure-focused network to effectively handle deghosting.
Sim-to-real domain adaptation. Domain adaptation has been widely used to transfer models trained on synthetic data to real-world settings. To address the domain shifts, researchers use either adversarial approaches (Ganin et al., 2016; Tzeng et al., 2017) or domain randomization (Tobin et al., 2017). Recently, adapter-based domain adaptation (Hu et al., 2021; Chen et al., 2022; Sung et al., 2022) has been proven to be more effective. Adapters (Hu et al., 2021; Chen et al., 2022) are a form of parameter-efficient fine-tuning (PEFT) (Hu et al., 2021; Zaken et al., 2021; Gao et al., 2021; Hu et al., 2022a), which require fewer parameters than full retraining and help mitigate catastrophic forgetting (Chen et al., 2022; Liu et al., 2023a) in domain adaptation. Additionally, Test-Time Adaptation (TTA) (Kundu et al., 2020; Boudiaf et al., 2023; Wang et al., 2022; Liu et al., 2023a; Chen et al., 2022) has been extensively explored, aiming to adapt a pre-trained model to unknown target domains during test-time, without any labeled or source domain data.
@@ 3 S2R-HDR Dataset
Previously, to create an HDR dataset with ground truth, researchers often use a beam splitter and two cameras to simultaneously capture images with two different exposures (Froehlich et al., 2014; Wang et al., 2021). The beam splitter only has two different exposures, which limits the dynamic range of the image. However, there are various high dynamic range scenarios in natural scenes, such as environments with direct sunlight. Accurately extracting tens of thousands of data samples from these scenes is a significant challenge. Previously, the largest commonly used dataset contained only 144 images (Tel et al., 2023), whereas ours includes 24,000 HDR images, representing a substantial leap in scale and diversity.
Moreover, capturing the ground truth often requires capturing different exposure images frame-by-frame (Kalantari and Ramamoorthi, 2017; Tel et al., 2023; Kong et al., 2024; Chen et al., 2021; Shu et al., 2024) and manually controlling motion between frames, making capturing extremely time-consuming. The captured motions are often limited and unrealistic, most of them are just basic human movements. These limitations have made it difficult to scale HDR datasets both in terms of size and motion variety. Below, we discuss how we solve all these challenges.
@@ 3.1 Rendering Design
Rendering high-quality HDR data presents several challenges. One challenge is that rendered images have a different distribution compared to the actual raw sensor data captured by cameras. To mitigate this difference, we made several improvements. First, by default, rendered images have a baked-in tone mapping, an irreversible process that compresses dynamic range for standard displays, making it hard to recover original HDR data. To overcome this, we design a custom UE5 (Unreal Engine 5) rendering pipeline that modifies tone mapping and gamma correction, ensuring the output remains in linear HDR space, and stores results in floating-point formats (EXR) to prevent data quantization. This approach ensures greater accuracy and makes the rendered data more suitable for HDR-related tasks. Second, we also simulate imperfections during handheld capturing. We incorporated camera shake simulation into our camera pose control to replicate the vibrations and instabilities that occur during real-world capture. This ensures that the rendered data closely mimics real-world shooting conditions, yielding more realistic HDR data for image processing and model training.
Another challenge is to construct realistic and diversified HDR scenes, with varying motion, lighting, and environmental details. To tackle this, we design and curate a diverse range of dynamic scene materials, including common moving objects such as animals, pedestrians, and vehicles, ensuring that the scenes exhibit a high degree of dynamism and complexity, as shown in Figure 2. Additionally, we carefully build a variety of high dynamic range scenes, encompassing both indoor and outdoor environments, various lighting conditions across different times of day, and extreme lighting scenarios. This diversity ensures that the generated HDR data simulates a broad range of real-world environments as much as possible. Additional examples of our motion materials and HDR scenes can be found in Appendix B.1 and Appendix B.2.
In total, we rendered 1,000 sequences, each containing 24 frames, resulting in a dataset of 24,000 HDR images, all stored in EXR format at a resolution of 1920 × 1080. As demonstrated in Figure 2, our rendered data encompasses a variety of environments and includes a broad range of motion types, showcasing a high degree of variability. Furthermore, since the data is in linear HDR format, it facilitates flexible data augmentation, enabling the easy generation of different LDR (low dynamic range) images, as shown in Figure 3.
@@ 3.2 Statistics and Analysis
Dataset FHLP $\uparrow$ EHL $\uparrow$ SI $\uparrow$ CF $\uparrow$ stdL $\uparrow$ ALL $\uparrow$ DR $\uparrow$ Size
Kalantari (Kalantari and Ramamoorthi, 2017) 15.07 3.07 18.4 4.74 10.02 6.19 2.71 89
SCT (Tel et al., 2023) 12.43 2.43 18.25 3.92 9.39 5.44 2.55 144
Challenge123 (Kong et al., 2024) 26.91 5.19 20.47 5.19 12.73 9.88 2.36 123
S2R-HDR 28.02 5.47 38.02 14.96 15.16 10.53 3.86 24000
We further analyze the diversity of S2R-HDR in comparison to previous datasets (Kalantari and Ramamoorthi, 2017; Tel et al., 2023; Kong et al., 2024). Following the methodology of Shu et al. (2024); Guo et al. (2023); Hu et al. (2022b), we use seven metrics to evaluate the diversity of different datasets across three dimensions: the extent of HDR, intra-frame diversity, and overall HDR style. As shown in Table 1, the S2R-HDR dataset outperforms all prior datasets across these metrics. The “Extent of HDR” metric demonstrates that our dataset covers a broader range of highlights, indicating an extended highlight range. The “Intra-frame Diversity” metric suggests that our images contain more detailed information and richer content. Finally, the “Overall Style” metric reveals that S2R-HDR exhibits a significantly higher dynamic range, surpassing the performance of previous datasets. Details of seven metrics can be found in Appendix  B.4.
Additionally, to visually illustrate the distribution between our dataset and existing real-world datasets (Kalantari and Ramamoorthi, 2017; Tel et al., 2023; Kong et al., 2024), we extract seven-dimensional feature vectors for each image and apply t-SNE (Van der Maaten and Hinton, 2008) for dimensionality reduction. As shown in Figure 8 (detailed in Appendix A.2), our S2R-HDR dataset spans a broader range in terms of data diversity. Additional data samples, along with optical flow, depth, and normal maps, are provided in Appendix B.6 and Appendix B.7, where we also discuss further application scenarios.
@@ 4 Domain Adaption
Figure 4: Structure of S2R-Adapter and t-SNE visualization of feature representations.
With all the careful design proposed in the previous section, there is still a noticeable gap between the synthetic S2R-HDR dataset and the real one, as shown in the t-SNE visualization in Appendix A.2. Thus, it is crucial to adapt the model trained on a large-scale rendered dataset to a small-scale real one. Still, direct fine-tuning on labeled real data can lead to overfitting and knowledge forgetting (Yosinski et al., 2014; Kirkpatrick et al., 2017).
@@
This is inspired by recent studies (Liu et al., 2023a; Chen et al., 2022; Sung et al., 2022), which suggest that adapters (Hu et al., 2021; Rebuffi et al., 2017; Chen et al., 2022) can mitigate forgetting in high-level vision tasks. Our adapter consists of two branches: a share branch to preserve shared knowledge from the rendered dataset, and a transfer branch to learn domain-specific knowledge from the real dataset, as shown in Figure 4 (a). We chose this design because we want to utilize both the shared knowledge from S2R-HDR to address large motion and dynamic range fusion, and the domain-specific knowledge from the real dataset, like more realistic textures.
More specifically, the proposed S2R-Adapter uses a plug-and-play structure, which can be attached to any pre-trained layers performing matrix multiplication (e.g., Linear Layer, Convolution Layer). Following  Liu et al. (2023a), we use a low-rank adapter as the share branch, which can better address knowledge forgetting, and use a high-rank adapter as the transfer branch, which can better extract domain-specific knowledge. Below we introduce details of each branch.
@@
$f=W_{0}x+\alpha_{s}\times f_{s}+\alpha_{t}\times f_{t}.\vskip-2.0pt$ (1)
@@
Verification using t-SNE. To verify the effectiveness of the proposed share branch and transfer branch adapters, we visualize the distributions of the rendered and real images using t-SNE (Van der Maaten and Hinton, 2008) in Figure 4 (b). From the share branch adapter, the feature distributions are consistent between the real and rendered domain, indicating that the share branch can ignore the domain difference between the real and the rendered domain, preserving the shared knowledge from forgetting. On the other hand, the transfer branch better separates the real distribution from the rendered distribution, showing its capability to model the real data distribution better and extract domain-specific knowledge in the real domain.
@@
Test-time adaptation with unlabeled data. During test-time adaptation, no labeled real domain data is available, and each sample is seen only once. Therefore, $\alpha_{s}$ and $\alpha_{t}$ cannot be learned across the real domain. Moreover, each test sample’s varying distance to the rendered domain requires adaptive scaling of transfer and shared branches. Therefore, inspired by Liu et al. (2023a); Ovadia et al. (2019), we dynamically adjust the scale factors using domain shift. For larger shifts, we increase the transfer branch’s scale factor, encouraging more knowledge from the real domain. For smaller shifts, we allocate more from the shared branch, preserving rendered domain knowledge. Domain shift is measured by model uncertainty, following Wang et al. (2022); Liu et al. (2023a); Roy et al. (2022); Ovadia et al. (2019). In our HDR Fusion task, we augment input samples $N$ times and calculate variance across $N$ outputs as the uncertainty value $\mathcal{U}(x)$. Augmentations include adjusting exposure, white balance, noise levels, and random flips. With the uncertainty value, we adaptively adjust scale factors:
$\alpha_{s}=1-\mathcal{U}(x);\quad\alpha_{t}=1+\mathcal{U}(x).\vskip-8.0pt$ (2)
Following previous works on test-time adaptation (Yang et al., 2024; Wang et al., 2022), we utilize the mean-teacher framework. As shown in Figure 5, we inject S2R-Adapters to both the teacher model $\mathcal{T}$ and the student model $\mathcal{S}$. We initialized both models with pre-trained weights on the rendered domain. Following Liu et al. (2023a), the teacher model generates uncertainty values and pseudo-labels $\widetilde{y}$ for updating the S2R-Adapters. The student model is optimized by the loss between the student output $\hat{y}$ and the pseudo-label $\widetilde{y}$. The teacher model updates via the exponential moving average (EMA) of the student model:
$\mathcal{T}^{t}=\lambda\mathcal{T}^{t-1}+(1-\lambda)\mathcal{S}^{t},\vskip-3.0pt$ (3)
where $t$ is the test step, $\lambda$ is set to $0.999$, following Tarvainen and Valpola (2017).